OpenStudy (anonymous):

can someone explain how to do this problem -6-n=2 n= or (there is a absolute sign around -6-n)

OpenStudy (lgbasallote):

\[| \, -6 - n \, | = 2\] is that the question?

OpenStudy (anonymous):

ya

OpenStudy (lgbasallote):

you just solve for n in -6 - n = 2 AND -6 - n = -2

OpenStudy (lgbasallote):

when -6 - n = 2 -6 -2 = n -8 = n

OpenStudy (lgbasallote):

when -6 - n = -2 -6 + 2 = n -4 = n

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

ya kind of i just dont get how you know how to add or subtract or whatever you have to do

OpenStudy (lgbasallote):

what do you mean i know how to add or subtract?

OpenStudy (anonymous):

you did -6+2=n

OpenStudy (lgbasallote):

-6 - n = -2 add n to both sides -6 = -2 + n add 2 to both sides -6 + 2 = n

OpenStudy (anonymous):

so when you add n it makes the number negative

OpenStudy (lgbasallote):

what becomes negative?

OpenStudy (anonymous):

the number when you add n

OpenStudy (lgbasallote):

nothing became engative

OpenStudy (lgbasallote):

i just added n

OpenStudy (lgbasallote):

-6 - n = -2 add n -6 - n + n = -2 + n -6 + 0 = -2 + n -6 = -2 + n nothing changed...n just became positive and went to the other side

OpenStudy (anonymous):

oh thanks alot but one last question why do you you have to do this at all

OpenStudy (lgbasallote):

you mean do math? i have no idea. i question that too

OpenStudy (anonymous):

lol